Find the length of the diagonal of a rectangle with a width of 16 inches and a length of 12 inches?
Answer by fractalier(6550)

Use the Pythagorean Theorem...the diagonal is the hypotenuse...
c^2 = a^2 + b^2
c^2 = 16^2 + 12^2
c^2 = 256 + 144
c^2 = 400
c = 20 inches